# Learning Decision-Focused Uncertainty Sets in Robust Optimization

### v4

@@ 3 Learning the uncertainty set
3.1 Constrained learning formulation
Consider a RO problem (1). We have a dataset of $N$ independent samples of the uncertain parameter $u$, $U^{N}=\{u^{i}\}_{i=1}^{N}\subseteq\mathcal{D}_{u}$, governed by the product distribution ${\mathbf{P}}^{N}_{u}$, and a dataset of $J$ independent samples of the family parameter $y$, $Y^{J}=\{y^{j}\}_{j=1}^{J}\subseteq\mathcal{D}_{y}$, governed by the product distribution ${\mathbf{P}}_{y}^{J}$. We define the combined samples $(u^{i},y^{j})$ for $i=1,\dots,N$ and $j=1,\dots,J$, and the combined dataset $U^{N}\times Y^{J},$ with product distribution defined as ${\mathbf{P}}^{N\times J}$. Our main requirement is to identify a set parameter $\theta$ for which the solutions $x(\theta,y)$ to problem (1) with uncertainty set $\mathcal{U}(\theta)$, satisfy the following finite-sample probabilistic guarantee
${\mathbf{P}}^{N\times J}\left\{{\mathbf{P}}_{(u,y)}(g(x(\theta,y),u,y)\leq 0)\geq 1-\eta\right\}\geq 1-\beta,$ (7)
where $\beta>0$ is the target confidence level across draws of the dataset $U^{N}\times Y^{J}$. Instead of working with the probability ${\mathbf{P}}_{(u,y)}$, which often leads to intractable formulations (Uryasev and Rockafellar,, 2001) as discussed in Section 2, we adopt a tractable approximation based on the conditional value at risk (Uryasev and Rockafellar,, 2001; Rockafellar and Uryasev,, 2002) defined in (3), obtaining the constraint
@@ 3.1 Constrained learning formulation
$\mathop{\bf CVaR}(g(x(\theta,y),u,y),\eta)\leq\kappa\quad\iff\quad\inf_{\alpha}\left\{\frac{{\mbox{\bf E}}_{(u,y)}(g(x(\theta,y),u,y)-\alpha)_{+}}{\eta}+\alpha\right\}\leq\kappa,$
where $\kappa>0$ is a threshold value that we define in Section 4 to obtain the finite sample guarantees (7). Our goal is to find the set parameter $\theta$ that minimizes the average objective of the RO problem across realizations of the family parameter $y$, which can be written as
@@
$\begin{array}[]{ll}\mbox{minimize}&{\mbox{\bf E}}_{y}\left(f(x(\theta,y))\right)\\ \mbox{subject to}&{\mbox{\bf E}}_{(u,y)}\left((g(x(\theta,y),u,y)-\alpha)_{+}/\eta+\alpha\right)\leq\kappa,\\ \end{array}$ (8)
@@
where $x(\theta,y)$ are solutions of the inner RO problems, defined as in (1). We given in Appendix A examples of parameterized uncertainty sets and the corresponding convex reformulations of (1), which we use to solve for $x(\theta,y)$. The expectation constraint corresponds to the $\mathop{\bf CVaR}$ constraint where $\alpha$ is shifted to become a minimization variable. When the functions $f$ and $g$ are convex in $x$, this shift in $\alpha$ has no impact on the optimal solution (Sarykalin et al.,, 2008). With this formulation, we tune $\alpha$ and the set parameter $\theta$ to improve the quality of the solutions $x(\theta,y)$ of the inner RO problem (1). Solving the training problem (8) is challenging because of the stochasticity in the objective and the constraints, and the nonconvex bi-level structure introduced by the inner RO problem. We address these challenges using a stochastic augmented Lagrangian algorithm.
@@ 3.2 Training the uncertainty set using a Stochastic augmented Lagrangian algorithm
As the distributions ${\mathbf{P}}_{u}$ and ${\mathbf{P}}_{y}$ are unknown, we are unable to obtain the true expectation in (8). As such, we work with sampled versions of the expectations. We define a condition on the sampled $\mathop{\bf CVaR}$,
$\widehat{\mathop{\bf CVaR}}(\theta)=\inf_{\alpha}\frac{1}{NJ}\sum_{j=1}^{J}\sum_{i=1}^{N}\left(\frac{(g(x(\theta,y^{j}),u^{i},y^{j})-\alpha)_{+}}{\eta}+\alpha\right)\leq\kappa,$ (9)
and require this condition to hold for feasible $\theta$. To find such a $\theta$, we introduce finite-sample approximations of the objective and constraint functions of problem (8),
$F(\theta)=\frac{1}{J}\sum_{j=1}^{J}f(x(\theta,y^{j}))\quad\text{and}\quad H(\alpha,\theta)=\frac{1}{NJ}\sum_{j=1}^{J}\sum_{i=1}^{N}(g(x(\theta,y^{j}),u^{i},y^{j})-\alpha)_{+}/\eta+\alpha,$
evaluated over the dataset $U^{N}\times Y^{J}$. We use these approximations to define the sampled version of problem (8),
$\begin{array}[]{ll}\mbox{minimize}&F(\theta)\\ \mbox{subject to}&H(\alpha,\theta)+s=\kappa\\ &s\geq 0,\\ \end{array}$ (10)
where $s$ is the non-negative slack variable in ${\mbox{\bf R}}^{d}$ to represent the inequality constraint as an equality. We can now write the augmented Lagrangian function as
$L(\alpha,s,\theta,\lambda,\mu)=F(\theta)$ $+\lambda(H(\alpha,\theta)+s-\kappa)+\frac{\mu}{2}\|H(\alpha,\theta)+s-\kappa\|^{2},$ (11)
where $\lambda$ is the Lagrange multiplier in R, and $\mu$ is the penalty parameter in R. This function $L$ is evaluated over the entire dataset $U^{N}\times Y^{J}$. We denote the estimated augmented Lagrangian over a subset (batch) of data $B=(U\times Y)\subseteq(U^{N}\times Y^{J})$ as $\hat{L}$. Using the stochastic augmented Lagrangian approach outlined in Algorithms 1 and 2, and under mild assumptions, we can find a feasible $\theta$ satisfying condition (9). We also define a slightly weaker optimality condition: an $\epsilon$-stationary point for problem (10). Given $\epsilon>0$, a point $(\alpha,s,\theta)\in{\mbox{\bf R}}^{q+2}$ is called an $\epsilon$-stationary point of (10) if there is a scalar $\lambda\in{\mbox{\bf R}}$ such that
$\|H(\alpha,\theta)+s-\kappa\|_{2}\leq\epsilon,\quad\mathop{\bf dist{}}(0,J_{F}(\theta)+\lambda J_{H+s-\kappa}(\alpha,s,\theta)))\leq\epsilon,$ (12)
where $J_{F}(\theta)$ is the conservative Jacobian (Definition 3.1 below) of the objective function $F$ in the variable $\theta$, $J_{H+s-\kappa}(\alpha,s,\theta)$ denotes the conservative Jacobian of the constraint function in variables $(\alpha,s,\theta)$, and the distance between a vector $x$ and set $\mathcal{X}$ is defined as $\mathop{\bf dist{}}{(x,\mathcal{X})}=\min_{y\in\mathcal{X}}\|x-y\|_{2}$. In addition, given $\lambda$ and $\mu$, a point $(\alpha,s,\theta)\in{\mbox{\bf R}}^{q+2}$ is called a stationary point of the augmented Lagrangian function (11), if $0\in J_{L}(\alpha,s,\theta)$. While useful in our convergence analysis, this stationarity condition is hard to evaluate. That’s why in practical applications, instead of using it to terminate Algorithm 2, we set the number of iterations $t_{\text{max}}$ to a reasonably large value.
@@
1:given $\theta^{0}=(A^{\text{init}},b^{\text{init}})$, $\alpha^{0}$, $s^{0}$, $\lambda^{0},\mu^{0},\epsilon\geq 0,\kappa$, $\sigma>1$, $\tau<1$, $H_{\text{best}}=\infty$
@@
3:     $(\alpha^{k},s^{k},\theta^{k})\leftarrow$ minimize $L(\alpha,s,\theta,\lambda^{k-1},\mu^{k-1})$ over $(\alpha,s,\theta)$ via Algorithm 2
4:     if $\|H(\alpha^{k},\theta^{k})+s^{k}-\kappa\|_{2}\leq\epsilon$ then
@@
6:     else if $\|H(\alpha^{k},\theta^{k})+s^{k}-\kappa\|_{2}\leq\tau\|H_{\text{best}}\|_{2}$ then
7:         $\lambda^{k}\leftarrow\Pi_{[{\lambda^{\rm{min}}},\lambda^{\rm{max}}]}(\lambda^{k-1}+\mu^{k-1}(H(\alpha^{k},\theta^{k})+s^{k}-\kappa)$) $\triangleright$ update Lagrange multipliers
8:         $H_{\text{best}}\leftarrow\|H(\alpha^{k},\theta^{k})+s^{k}-\kappa\|_{2}$
@@
1:given $\theta^{0}$, $\alpha^{0}$, $s^{0}$, $\lambda,\mu,\{\delta^{t}\}_{t\in\mathbf{N}},U^{N}\times Y^{J}$
@@
3:     sample data $(U^{t}\times Y^{t})\subseteq(U^{N}\times Y^{J})$
4:     $x^{t}(\theta^{t-1},y)\leftarrow$ solve inner RO problem (1) for all $y\in Y^{t}$
5:     $J_{\hat{L}}\leftarrow J_{\hat{L}}(\alpha^{t-1},\theta^{t-1},s^{t-1})$triangleright compute Jacobian of $\hat{L}$ over $U^{t}\times Y^{t}$
@@
7:     $s^{t}\leftarrow(s^{t-1}-\delta^{t}v_{s}^{t})_{+},\quad\text{with}\quad v_{s}^{t}\in J_{\hat{L}}(s)$
8:     $\theta^{t}\leftarrow\theta^{t-1}-\delta^{t}v_{\theta}^{t},\quad\text{with}\quad v_{\theta}^{t}\in J_{\hat{L}}(\theta)$
9:return $(\alpha^{t},s^{t},\theta^{t})$
@@ 3.2.1 Path differentiability and conservative Jacobians
Given a locally Lipschitz continuous function $F:{\mbox{\bf R}}^{m}\rightarrow{\mbox{\bf R}}^{m}$, we call $J_{F}:{\mbox{\bf R}}^{n}\rightrightarrows{\mbox{\bf R}}^{n\times m}$ its conservative Jacobian if $J_{F}$ has a closed graph, is locally bounded, and is nonempty with
@@
for any absolutely continuous curve $\gamma$ in ${\mbox{\bf R}}^{n}$. A locally Lipschitz function $F$ that admits a conservative Jacobian $J_{F}$ is called path differentiable.
@@
The Lagrangian function $L(\alpha,s,\theta,\lambda,\mu)$ is locally Lipschitz in variables $\alpha,s,\theta,\lambda,\mu$. A function $F:{\mbox{\bf R}}^{n}\rightarrow{\mbox{\bf R}}^{m}$ is locally Lipschitz if, for each $x\in{\mbox{\bf R}}^{n}$, there exists a neighborhood $\mathcal{V}$ of $x$ such that $F$ is Lipschitz on $\mathcal{V}$.
We also require the augmented Lagrangian to be locally Lipschitz continuous, as definable locally Lipschitz functions are shown to be path differentiable, meaning they admit conservative Jacobians and follow the chain rule (Bolte et al.,, 2021). Admittance of the chain rule is imperative, as $L$ depends on $\theta$ indirectly through the solution $x(\theta,y)$ of the inner robust problem (1). In order to establish the conservative Jacobians of $L$ in the variable $\theta$, we thus need to show the path differentiability of the inner problem and compute $J_{x}(\theta)$, for which we require a uniqueness assumption.
@@
Suppose Assumption 3.3 holds. Let $x(\theta,y)$ be the unique solution, for a given input pair $(\theta,y)$, to the convex conic reformulation of the inner optimization problem (1). This solution is path differentiable with nonempty conservative Jacobian $J_{x}(\theta)\neq\emptyset$.
We can then evaluate the conservative Jacobians of the augmented Lagrangian function.
@@
Suppose Assumptions 3.1 and 3.2 hold, and for each given input pair $(\theta,y)$, the solution $x(\theta,y)$ of the optimization problem (1) is path differentiable with conservative Jacobian $J_{x}(\theta)$. Then, the Lagrangian function $L(\alpha,s,\theta,\lambda,\mu)$, from equation (11), and its batched version $\hat{L}(\alpha,s,\theta,\lambda,\mu)$, are path differentiable with conservative Jacobians $J_{L}(\alpha,s,\theta)$ and $J_{\hat{L}}(\alpha,s,\theta)$.
The specific forms of the conservative Jacobians established in Proposition 3.1 are given in Appendix B.
@@ 3.2.2 Convergence analysis
There exists a constant $\alpha^{\rm max}>0$ and compact set $\mathcal{D}_{\alpha}\subset{\mbox{\bf R}}$, such that, for $\alpha^{0}\in\mathcal{D}_{\alpha}$, $\|\alpha^{t}\|\leq\alpha^{\rm max}$ almost surely. Similarly, there exist constants $\theta^{\rm max},s^{\rm max},x^{\rm max},y^{\rm max}>0$, compact sets $\mathcal{D}_{\theta}\subset{\mbox{\bf R}}^{q}$, $\mathcal{D}_{s}\subset{\mbox{\bf R}}$, and $\mathcal{X}\subset{\mbox{\bf R}}^{n}$, such that, for $\theta^{0}\in\mathcal{D}_{\theta}$, $s^{0}\in\mathcal{D}_{s}$, and $y\in\mathcal{D}_{y}$, we have $\|\theta^{t}\|\leq\theta^{\rm max}$, $\|s^{t}\|\leq s^{\rm max}$, and $x(\theta^{t},y)\in\mathcal{X}=\{x\in{\mbox{\bf R}}^{n}\mid\|x\|\leq x^{\rm max}\}$ almost surely. In addition, the optimal Lagrangian multiplier $\lambda^{\star}\in[\lambda^{\rm{min}}$, $\lambda^{\rm{max}}]$.
@@
Under these assumptions, by Theorem 6.2 of Davis et al., (2018) and Theorem 3 of Bolte et al., (2021), the sequence $\{\alpha^{t},s^{t},\theta^{t}\}$ in Algorithm 2 converges to a stationary point of the Lagrangian function (11) with arbitrary $\lambda$ and $\mu$.
We note that $\lambda^{k}$ and $\mu^{k}$ in Algorithm 1 are potentially unbounded, in which case they do not converge. The stationary point $(\alpha^{k},s^{k},\theta^{k})$ of the augmented Lagrangian function for any $\lambda^{k}$ and $\mu^{k}$ is, therefore, not guaranteed to be a feasible solution of the $\mathop{\bf CVaR}$-constrained problem (10). However, with the use of the safeguards $[\lambda^{\rm{min}}$, $\lambda^{\rm{max}}]$, we can guarantee that the limit point of the augmented Lagrangian algorithm is at least a stationary point of constraint violation (Birgin and Martinez,, 2009, Theorem 6.3). More specifically, below we show that the algorithm finds a feasible point of condition (9) under mild assumptions, and otherwise converges to a stationary point of the constraint violation function $(1/2)\|H(\alpha,\theta)+s-\kappa\|^{2}$. The proof is delayed to Appendix D.
@@
Suppose Assumptions 3.1 to 3.5 hold. If the sequence of penalty parameters $\{\mu^{k}\}$ is bounded, then the sequence of updates $\{\theta^{k}\}$ in Algorithm 1 converges almost surely to a feasible point of the $\widehat{\mathop{\bf CVaR}}$ condition (9) as $k\to\infty$. Otherwise, $\{\alpha^{k},s^{k},\theta^{k}\}$ converges to a stationary point of the constraint violation function $(1/2)\|H(\alpha,\theta)+s-\kappa\|^{2}$.
Note that a feasible point $\theta^{\star}$ of the $\widehat{\mathop{\bf CVaR}}$ condition (9) also implies a feasible point of the contrained problem (10). In practice, with a finite number of iterations and some $\epsilon>0$, we obtain an $\epsilon$-stationary solution of (10) satisfying condition (12).

### v5

@@ 3 Learning the uncertainty set
3.1 Constrained stochastic learning formulation
As the distribution ${\mathbf{P}}_{(u,x)}$ is unknown, we are unable to obtain the true expectation in (3). We do, however, assume access to historical data. Suppose we are given a pair of datasets $W^{N}=(U^{N},X^{N})=\{(u^{i},x^{i})\}_{i=1}^{N}\subseteq\mathcal{D}_{(u,x)}$ of a set of $N$ independent samples of the uncertain parameter $u$ and context parameter $x$, governed by ${\mathbf{P}}^{N}$, their joint distribution. As the true probabilistic guarantee (2) cannot be obtained, we instead target the the following finite-sample probabilistic guarantee
${\mathbf{P}}^{N}\left({\mathbf{P}}_{(u,x)}(g(z(\theta,x),u,x)\leq 0)\geq 1-\eta\right)\geq 1-\beta,$ (11)
where $\beta>0$ is the target confidence level across draws of the dataset $W^{N}$. In other words, we are looking for $\theta$ such that the probability of observing $N$ samples which results in data-driven decisions $z(\theta,x)$ that do not satisfy the desired guarantee is small. To obtain such a guarantee, we first ensure the convergence of a data-driven version of (3). To this end, we define a constraint on the sampled $\mathop{\bf CVaR}$,
@@ 3.1 Constrained stochastic learning formulation
$\widehat{\mathop{\bf CVaR}}(\theta)=\inf_{\alpha}\frac{1}{N}\sum_{i=1}^{N}\left(\frac{(g(z(\theta,x^{i}),u^{i},x^{i})-\alpha)_{+}}{\eta}+\alpha\right)\leq\kappa,$ (12)
and require this condition to hold for feasible $\theta$. The value $\kappa<0$ is a threshold margin that we define later in Section 4, to ensure the finite sample guarantee (11). We can then formulate the stochastic objective and constraint functions as follows,
$F(\theta)=\frac{1}{N}\sum_{i=1}^{N}\Phi(z(\theta,x^{i}),u^{i},x^{i})\quad\text{and}\quad H(\alpha,\theta)=\frac{1}{N}\sum_{i=1}^{N}((g(z(\theta,x^{i}),u^{i},x^{i})-\alpha)_{+}/\eta+\alpha-\kappa)_{+},$
evaluated over the dataset $W^{N}$. We use these approximations to define the sampled version of (3),
@@
$\begin{array}[]{ll}\underset{\alpha,\theta}{\mbox{minimize}}&F(\theta)\\ \mbox{subject to}&H(\alpha,\theta)=0,\\ \end{array}$ (13)
@@
where the constraint $H$ corresponds to the $\mathop{\bf CVaR}$ constraint where $\alpha$ is also a minimization variable. When the functions $f$ and $g$ are convex in $z$, optimizing over $\alpha$ as well has no impact on the optimal solution (Sarykalin et al.,, 2008). With this formulation, we thus tune $\alpha$ and the set parameter $\theta$ to improve the quality of the solutions $z(\theta,x^{i})$ of the inner RO problem (1), measured by the evaluation metric $\Phi(\cdot)$. We give in Appendix A examples of parameterized uncertainty sets and the corresponding convex reformulations of (1), which we use to solve for $z(\theta,x^{i})$. In the motivating example above, we gave a parametrization of an ellipsoidal uncertainty set (10), where the trainable parameters are $(W_{A},W_{b},h_{A},h_{b})$. Learning $\theta$ using (13) thus equates to learning these parameters, i.e., learning a feasible and ideally optimal mapping from context parameters to a region of uncertainty realizations.
Solving the training problem (13) is challenging because of the nonconvex bi-level structure introduced by the inner RO problem. While it may be possible to directly reformulate it as a single-level problem, this would involve several bilinear constraints over the entire training dataset, which can be intractable. We address these challenges in the following section, by introducing a stochastic first-order method to address problem (13) with tractable iterations.
@@ 3.2 Training the uncertainty set using a stochastic augmented Lagrangian algorithm
To solve (13), we construct the augmented Lagrangian function
$L(\alpha,\theta,\lambda,\mu)=F(\theta)$ $+\lambda H(\alpha,\theta)+\frac{\mu}{2}\|H(\alpha,\theta)\|^{2},$ (14)
where $\lambda$ is the Lagrange multiplier in ${\mathbf{R}}$, and $\mu$ is the penalty parameter in ${\mathbf{R}}$. This function $L$ is evaluated over the entire dataset $W^{N}$. We denote the estimated augmented Lagrangian over a subset (batch) of data $B\subseteq W^{N}$ as $\hat{L}$. Using the stochastic augmented Lagrangian approach outlined in Algorithms 1 and 2, and under mild assumptions, we can find a feasible $\theta$ satisfying condition (12). We also define a slightly weaker optimality condition: an $\epsilon$-stationary point for problem (13). Given $\epsilon>0$, a point $(\alpha,\theta)\in{\mathbf{R}}^{q+1}$ is called an $\epsilon$-stationary point of (13) if there is a scalar $\lambda\in{\mathbf{R}}$ such that
$\|H(\alpha,\theta)\|_{2}\leq\epsilon,\quad\mathop{\bf dist{}}(0,J_{F}(\theta)+\lambda J_{H}(\alpha,\theta)))\leq\epsilon,$ (15)
where $J_{F}(\theta)$ is the conservative Jacobian (Definition 3.1 below) of the objective function $F$ in the variable $\theta$, $J_{H}(\alpha,\theta)$ denotes the conservative Jacobian of the constraint function in variables $(\alpha,\theta)$, and the distance between a vector $z$ and set $\mathcal{Z}$ is defined as $\mathop{\bf dist{}}{(z,\mathcal{Z})}=\min_{y\in\mathcal{Z}}\|z-y\|_{2}$. In addition, given $\lambda$ and $\mu$, a point $(\alpha,\theta)\in{\mathbf{R}}^{q+1}$ is called a stationary point of the augmented Lagrangian function (14), if $0\in J_{L}(\alpha,\theta)$.
@@
1:given $\theta^{0}=(A^{\text{init}},b^{\text{init}})$, $\alpha^{0}$, $\lambda^{0},\mu^{0},\epsilon\geq 0,\kappa$, $\sigma>1$, $\tau<1$, $\eta\in(0,1)$, $H_{\text{best}}=\infty$
@@
3:     $(\alpha^{k},\theta^{k})\leftarrow$ minimize $L(\alpha,\theta,\lambda^{k-1},\mu^{k-1})$ over $(\alpha,\theta)$ via Algorithm 2
4:     if $\|H(\alpha^{k},\theta^{k})\|_{2}\leq\epsilon$ then
@@
6:     else if $\|H(\alpha^{k},\theta^{k})\|_{2}\leq\tau\|H_{\text{best}}\|_{2}$ then
7:         $\lambda^{k}\leftarrow\Pi_{[{\lambda^{\rm{min}}},\lambda^{\rm{max}}]}(\lambda^{k-1}+\mu^{k-1}(H\alpha^{k},\theta^{k})$) $\triangleright$ update Lagrange multipliers
8:         $H_{\text{best}}\leftarrow\|H(\alpha^{k},\theta^{k})\|_{2}$
@@
1:given $\theta^{0}$, $\alpha^{0}$, $\lambda,\mu,\{\delta^{t}\}_{t\in\mathbf{N}},W^{N}$
@@
3:     sample data $W^{t}=(U^{t},X^{t})\subseteq W^{N}$
4:     $z^{t}(\theta^{t-1},x)\leftarrow$ solve inner RO problem (1) for all $x\in X^{t}$
5:     $J_{\hat{L}}\leftarrow J_{\hat{L}}(\alpha^{t-1},\theta^{t-1})$triangleright compute Jacobian of $\hat{L}$ over $W^{t}$
@@
7:     $\theta^{t}\leftarrow\theta^{t-1}-\delta^{t}v_{\theta}^{t},\quad\text{with}\quad v_{\theta}^{t}\in J_{\hat{L}}(\theta)$
8:return $(\alpha^{t},s^{t},\theta^{t})$
@@ 3.2.1 Path differentiability and conservative Jacobians
Given a locally Lipschitz continuous function $F:{\mathbf{R}}^{m}\rightarrow{\mathbf{R}}^{m}$, we call $J_{F}:{\mathbf{R}}^{n}\rightrightarrows{\mathbf{R}}^{n\times m}$ its conservative Jacobian if $J_{F}$ has a closed graph, is locally bounded, and is nonempty with
@@
for any absolutely continuous curve $\gamma$ in ${\mathbf{R}}^{n}$. A locally Lipschitz function $F$ that admits a conservative Jacobian $J_{F}$ is called path differentiable.
@@
The Lagrangian function $L(\alpha,\theta,\lambda,\mu)$ is locally Lipschitz in variables $\alpha,\theta,\lambda,\mu$. A function $F:{\mathbf{R}}^{n}\rightarrow{\mathbf{R}}^{m}$ is locally Lipschitz if, for each $z\in{\mathbf{R}}^{n}$, there exists a neighborhood $\mathcal{V}$ of $z$ such that $F$ is Lipschitz on $\mathcal{V}$.
We also require the augmented Lagrangian to be locally Lipschitz continuous, as definable locally Lipschitz functions are shown to be path differentiable, meaning they admit conservative Jacobians and follow the chain rule (Bolte et al.,, 2021). Admittance of the chain rule is imperative, as $L$ depends on $\theta$ indirectly through the solution $z(\theta,x)$ of the inner robust problem (1). In order to establish the conservative Jacobians of $L$ in the variable $\theta$, we thus need to show the path differentiability of the inner problem and compute $J_{z}(\theta)$, for which we require a uniqueness assumption.
@@
Suppose Assumption 3.3 holds. Let $z(\theta,x)$ be the unique solution, for a given input pair $(\theta,x)$, to the convex conic reformulation of the inner optimization problem (1). This solution is path differentiable with nonempty conservative Jacobian $J_{z}(\theta)\neq\emptyset$.
We can then evaluate the conservative Jacobians of the augmented Lagrangian function. The specific forms of the conservative Jacobians established in Proposition 3.1 are given in Appendix B.
@@
Suppose Assumptions 3.1 and 3.2 hold, and for each given input pair $(\theta,x)$, the solution $z(\theta,x)$ of the optimization problem (1) is path differentiable with conservative Jacobian $J_{z}(\theta)$. Then, the Lagrangian function $L(\alpha,\theta,\lambda,\mu)$, from equation (14), and its batched version $\hat{L}(\alpha,\theta,\lambda,\mu)$, are path differentiable with conservative Jacobians $J_{L}(\alpha,\theta)$ and $J_{\hat{L}}(\alpha,\theta)$.
@@ 3.2.2 Convergence analysis
There exists a constant $\alpha^{\rm max}>0$ and compact set $\mathcal{D}_{\alpha}\subset{\mathbf{R}}$, such that for $\alpha^{0}\in\mathcal{D}_{\alpha}$, $\|\alpha^{t}\|\leq\alpha^{\rm max}$ almost surely. Similarly, there exist constants $\theta^{\rm max},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}z^{\rm max}},y^{\rm max}>0$, compact sets $\mathcal{D}_{\theta}\subset{\mathbf{R}}^{q}$, $\mathcal{Z}\subset{\mathbf{R}}^{n}$, such that for $\theta^{0}\in\mathcal{D}_{\theta}$ and $x\in\mathcal{D}_{x}$, we have $\|\theta^{t}\|\leq\theta^{\rm max}$ and $z(\theta^{t},x)\in\mathcal{Z}=\{z\in{\mathbf{R}}^{n}\mid\|x\|\leq z^{\rm max}\}$ almost surely. In addition, the optimal Lagrangian multiplier $\lambda^{\star}\in[\lambda^{\rm{min}}$, $\lambda^{\rm{max}}]$.
@@
Under these assumptions, by Theorem 6.2 of Davis et al., (2018) and Theorem 3 of Bolte et al., (2021), the sequence $\{\alpha^{t},\theta^{t}\}$ in Algorithm 2 converges to a stationary point of the Lagrangian function (14) with arbitrary $\lambda$ and $\mu$. We note that $\lambda^{k}$ and $\mu^{k}$ in Algorithm 1 are potentially unbounded, in which case they do not converge. The stationary point $(\alpha^{k},\theta^{k})$ of the augmented Lagrangian function for any $\lambda^{k}$ and $\mu^{k}$ is, therefore, not guaranteed to be a feasible solution of the $\mathop{\bf CVaR}$-constrained problem (13). However, with the use of the safeguards $[\lambda^{\rm{min}}$, $\lambda^{\rm{max}}]$, we can guarantee that the limit point of the augmented Lagrangian algorithm is at least a stationary point of constraint violation (Birgin and Martinez,, 2009, Theorem 6.3). More specifically, below we show that the algorithm finds a feasible point of condition (12) under mild assumptions, and otherwise converges to a stationary point of the constraint violation function $(1/2)\|H(\alpha,\theta)\|^{2}$. The proof is delayed to Appendix D.
@@
Suppose Assumptions 3.1 to 3.5 hold. If the sequence of penalty parameters $\{\mu^{k}\}$ is bounded, then the sequence of updates $\{\theta^{k}\}$ in Algorithm 1 converges almost surely to a feasible point, $\theta^{\star}$, of the $\widehat{\mathop{\bf CVaR}}$ condition (12) as $k\to\infty$. Otherwise, $\{\alpha^{k},\theta^{k}\}$ converges to a stationary point of the constraint violation function $(1/2)\|H(\alpha,\theta)\|^{2}$.
Note that a feasible point $\theta^{\star}$ of the $\widehat{\mathop{\bf CVaR}}$ condition (12) also implies a feasible point of the constrained problem (13). In practice, with a finite number of iterations and some $\epsilon>0$, we obtain an $\epsilon$-stationary solution of (13) satisfying condition (15).